# Dynamical Edge Modes in $p$-form Gauge Theories

### v1

@@ 4.5.2 $p$-form gauge theory
We now carry out a similar analysis for $p$-form gauge theory. For simplicity we assume $\Sigma_{\varepsilon}$ has the topology of a $(D-1)$-ball and that $22\leq p\leq D-$ 2. Once again we modify the PMC boundary condition by specifying a $\tau$-independent electric flux on the boundary. Now the equation of motion requires it to be co-closed, so we write it as $E^{T}_{\perp}$. The partition function defined by this boundary condition factorizes into classical and fluctuation parts,
@@
Here $|\mathcal{C}^{p-1}(\partial\Sigma_{\varepsilon})|$ is formally the (infinite) volume of the space of co-closed $(p-1)$-forms on $\mathcal{M}_{\varepsilon}$, and the formal quotient reduces to
$\frac{|\mathcal{C}^{p-1}(\partial\Sigma_{\varepsilon})|}{|\mathcal{G}^{{\partial\Sigma_{\varepsilon}}}|}=|U(1)|^{(-)^{p}}\prod_{k=0}^{p-2}\det^{\prime}\left(\frac{\Delta_{k}^{T,\partial\Sigma}}{\mu^{2}}\right)^{-\frac{(-)^{p-k}}{2}}\;.$ (4.86)
@@
^$\hat{\nabla}^{j}(\sqrt{g^{\tau\tau}}\gamma_{ji_{1}\dots i_{p-2}})=0\;,\qquad\frac{1}{\sqrt{g^{\tau\tau}}}\hat{\nabla}^{j}(\sqrt{g^{\tau\tau}}p!\hat{\nabla}_{[j}\gamma_{i_{1}\dots i_{p-1}]})=0\;.$ (4.88)
@@ Acknowledgments
It is a great pleasure to thank Changha Choi, Jaume Gomis, and Jacob McNamara for useful discussions. AB is supported by the Celestial Holography Initiative at the Perimeter Institute for Theoretical Physics and the Simons Collaboration on Celestial Holography. Research at the Perimeter Institute is supported by the Government of Canada through the Department of Innovation, Science and Industry Canada and by the Province of Ontario through the Ministry of Colleges and Universities. AL was supported by the Stanford Science Fellowship.
@@ Appendix A Conventions for differential forms
$(*_{M}\omega)_{\mu_{1}\dots\mu_{D-k}}\equiv\frac{1}{k!}\varepsilon_{\mu_{1}\dots\mu_{D-k}}{}^{\nu_{1}\dots\nu_{k}}\omega_{\nu_{1}\dots\nu_{k}}.$ (A.4)
@@ Appendix B Horizon limit calculation
00 0 $=g_{a_{1}b_{1}}\dots g_{a_{p-1}b_{p-1}}\frac{-p}{S\sqrt{\hat{g}}}\partial_{i}\left(S\sqrt{\hat{g}}\,g^{ij}g^{b_{1}c_{1}}\dots g^{b_{p-1}c_{p-1}}\partial_{[j}\beta_{c_{1}\dots c_{p-1}]}\right)$ (B.3)
@@
[…]
@@
00 0 $\approx\frac{-p}{S}\partial_{r}\left(S\partial_{[r}\beta_{a_{1}\dots a_{p-1}]}\right)+\left(\Delta^{\partial\Sigma}_{p-1}\beta\right)_{a_{1}\dots a_{p-1}}+(p-1)\partial_{[a_{1}}\left(g_{a_{2}|b_{2}|}\dots g_{a_{p-1}]b_{p-1}}\frac{1}{\sqrt{\hat{g}}}\partial_{c}\left(\sqrt{\hat{g}}\,\beta^{cb_{2}\dots b_{p-1}}\right)\right)$ (B.5)
@@
00 0 $\approx\frac{-p}{S}\partial_{r}\left(S\partial_{[r}\beta_{a_{1}\dots a_{p-1}]}\right)+\left(\Delta^{\partial\Sigma}_{p-1}\beta\right)_{a_{1}\dots a_{p-1}}-(p-1)\partial_{[a_{1}}\left(g_{a_{2}|b_{2}|}\dots g_{a_{p-1}]b_{p-1}}\frac{1}{S\sqrt{\hat{g}}}\partial_{r}\left(S\sqrt{\hat{g}}\,\beta^{rb_{2}\dots b_{p-1}}\right)\right)$ (B.6)
@@
00 0 $\approx\frac{-1}{S}\partial_{r}\left(S\partial_{r}\beta_{a_{1}\dots a_{p-1}}\right)+\frac{p-1}{S}\partial_{r}\left(S\partial_{[a_{1}}\beta_{|r|a_{2}\dots a_{p-1}]}\right)+\left(\Delta^{\partial\Sigma}_{p-1}\beta\right)_{a_{1}\dots a_{p-1}}-\frac{p-1}{S}\partial_{r}\left(S\partial_{[a_{1}}\beta_{|r|a_{2}\dots a_{p-1}]}\right)$ (B.7)
@@ Appendix C $dS_{D}$ static patch and $S^{D}$ partition functions with no brick wall
[…]
@@
From (C.2), we that that a massless $p$-form corresponds to $\Delta=D-1-p$ or $\Delta=p$. Some of the degrees of freedom become pure-gauge in the massless limit, which is reflected in the naïve character
@@ C.2 Laplacians for transverse $p$-forms on $S^{D}$
The eigenvalues and degeneracies of the Laplacian $\Delta^{T}_{p}\equiv dd^{\dagger}+d^{\dagger}d$ acting on a transverse $p$-form ($11\leq p\leq D-$ 1) on a round $S^{D}$ with radius $R$ are [89, 60, 73]
@@ C.3 Massive $p$-forms on $S^{D}$
and $d^{D}_{n,(p)}$ is the degeneracy (C.2) for the $n$-th transverse $p$-form spherical harmonics on $S^{D}$. Here we focus on the case of the principal series (C.1); the complementary case is obtained by analytic continuation. To derive (C.22) from (4.54), one decomposes the $p$-form into a transverse and longitudinal part: $A_{\mu_{1}\cdots\mu_{p}}=A^{T}_{\mu_{1}\cdots\mu_{p}}+p\nabla_{[\mu_{1}}C_{\mu_{2}\cdots\mu_{p}]}$. The integration over the transverse part $A^{T}$ leads to the $n\geq 1$ terms in (C.22). On the other hand, $C_{\mu_{1}\cdots\mu_{p-1}}$ only enters the action through the mass term in the action (4.54); upon integration, this leads to an infinite product of $m^{2}$. Absorbing this into a local counterterm in a way consistent with locality leads to the $n=-p$ or $n=p-(D-1)$ term in the range $0\geq n\geq-\lfloor\frac{D-1}{2}\rfloor$ in (C.22). This is done explicitly for $p=1$ in [90].
@@ C.4 Massless $p$-form gauge fields on $S^{D}$
$\mathcal{Z}^{p}_{\rm PI}\left[S^{D}\right]=\left(\frac{\mu^{p+1}\sqrt{2\pi R^{D}{\rm Vol}\left(S^{D}\right)}}{q}\right)^{(-)^{p}}\left(\det^{\prime}\frac{\Delta_{0}}{\mu^{2}}\right)^{-\frac{(-)^{p}}{2}}\prod_{k=1}^{p}\left(\det\frac{\Delta^{T}_{k}}{\mu^{2}}\right)^{-\frac{(-)^{p-k}}{2}}\;.$ (C.28)

### v2

@@ 4.5.2 $p$-form gauge theory
We now carry out a similar analysis for $p$-form gauge theory. For simplicity we assume $\Sigma_{\varepsilon}$ has the topology of a (D1)(D-1)-ball and that $2\leq p\leq D-2$. Once again we modify the PMC boundary condition by specifying a $\tau$-independent electric flux on the boundary. Now the equation of motion requires it to be co-closed, so we write it as $E^{T}_{\perp}$. The partition function defined by this boundary condition factorizes into classical and fluctuation parts,
@@
Here $|\mathcal{C}^{p-1}(\partial\Sigma_{\varepsilon})|$ is formally the (infinite) volume of the space of co-closed (p1)(p-1)-forms on $\mathcal{M}_{\varepsilon}$, and the formal quotient reduces to
$\frac{|\mathcal{C}^{p-1}(\partial\Sigma_{\varepsilon})|}{|\mathcal{G}^{{\partial\Sigma_{\varepsilon}}}|}=|U(1)|^{(-)^{p}}\prod_{k=0}^{p-2}\det\hskip 0.0pt^{\prime}\left(\frac{\Delta_{k}^{T,\partial\Sigma}}{\mu^{2}}\right)^{-\frac{(-)^{p-k}}{2}}\;.$ (4.86)
@@
$\hat{\nabla}^{j}(\sqrt{g^{\tau\tau}}\gamma_{ji_{1}\dots i_{p-2}})=0\;,\qquad\frac{1}{\sqrt{g^{\tau\tau}}}\hat{\nabla}^{j}(\sqrt{g^{\tau\tau}}p!\hat{\nabla}_{[j}\gamma_{i_{1}\dots i_{p-1}]})=0\;.$ (4.88)
@@ Acknowledgments
It is a great pleasure to thank Changha Choi, Jaume Gomis, and Jacob McNamara for useful discussions. AB is supported by the Celestial Holography Initiative at the Perimeter Institute for Theoretical Physics and the Simons Collaboration on Celestial Holography. Research at the Perimeter Institute is supported by the Government of Canada through the Department of Innovation, Science and Industry Canada and by the Province of Ontario through the Ministry of Colleges and Universities. AL was supported by the Stanford Science Fellowship and a Simons Investigator award.
@@ Appendix A Conventions for differential forms
$(*_{M}\omega)_{\mu_{1}\dots\mu_{D-k}}\equiv\frac{1}{k!}\,\omega_{\nu_{1}\dots\nu_{k}}\varepsilon^{\nu_{1}\dots\nu_{k}}{}_{\mu_{1}\dots\mu_{D-k}}$. (A.4)
@@ Appendix B Horizon limit calculation
$0$ $=g_{a_{1}b_{1}}\dots g_{a_{p-1}b_{p-1}}\frac{-p}{S\sqrt{\hat{g}}}\partial_{i}\left(S\sqrt{\hat{g}}\,g^{ij}g^{b_{1}c_{1}}\dots g^{b_{p-1}c_{p-1}}\partial_{[j}\beta_{c_{1}\dots c_{p-1}]}\right)$ (B.3)
@@
In the third line the use of $\approx\,$ indicates that we are dropping subleading terms like $\partial_{r}g_{ab}$ and $S^{-1}\partial_{a}S$. The latter term is closely related to the Laplacian on $\partial\Sigma$, acting on (p1)(p-1)-forms,
@@
$0$ $\approx\frac{-p}{S}\partial_{r}\left(S\partial_{[r}\beta_{a_{1}\dots a_{p-1}]}\right)+\left(\Delta^{\partial\Sigma}_{p-1}\beta\right)_{a_{1}\dots a_{p-1}}+(p-1)\partial_{[a_{1}}\left(g_{a_{2}|b_{2}|}\dots g_{a_{p-1}]b_{p-1}}\frac{1}{\sqrt{\hat{g}}}\partial_{c}\left(\sqrt{\hat{g}}\,\beta^{cb_{2}\dots b_{p-1}}\right)\right)$ (B.5)
@@
$0$ $\approx\frac{-p}{S}\partial_{r}\left(S\partial_{[r}\beta_{a_{1}\dots a_{p-1}]}\right)+\left(\Delta^{\partial\Sigma}_{p-1}\beta\right)_{a_{1}\dots a_{p-1}}-(p-1)\partial_{[a_{1}}\left(g_{a_{2}|b_{2}|}\dots g_{a_{p-1}]b_{p-1}}\frac{1}{S\sqrt{\hat{g}}}\partial_{r}\left(S\sqrt{\hat{g}}\,\beta^{rb_{2}\dots b_{p-1}}\right)\right)$ (B.6)
@@
$0$ $\approx\frac{-1}{S}\partial_{r}\left(S\partial_{r}\beta_{a_{1}\dots a_{p-1}}\right)+\frac{p-1}{S}\partial_{r}\left(S\partial_{[a_{1}}\beta_{|r|a_{2}\dots a_{p-1}]}\right)+\left(\Delta^{\partial\Sigma}_{p-1}\beta\right)_{a_{1}\dots a_{p-1}}-\frac{p-1}{S}\partial_{r}\left(S\partial_{[a_{1}}\beta_{|r|a_{2}\dots a_{p-1}]}\right)$ (B.7)
@@ Appendix C $dS_{D}$ static patch and $S^{D}$ partition functions with no brick wall
$\text{Principal series}:$ $\qquad\Delta=\frac{D-1}{2}+i\nu\;,\qquad\nu\in\mathbb{R}$
$\text{Complementary series}:$ $\qquad\Delta=\frac{D-1}{2}+\nu\;,\qquad|\nu|<\left|\frac{D-1}{2}-p\right|\;.$ (C.3)
@@
From (C.2), we see that a massless $p$-form corresponds to $\Delta=D-1-p$ or $\Delta=p$. Some of the degrees of freedom become pure-gauge in the massless limit, which is reflected in the naïve character
@@ C.2 Laplacians for transverse $p$-forms on $S^{D}$
The eigenvalues and degeneracies of the Laplacian $\Delta^{T}_{p}\equiv dd^{\dagger}+d^{\dagger}d$ acting on a transverse $p$-form ($1\leq p\leq D-1$) on a round $S^{D}$ with radius $R$ are [102, 63, 86]
@@ C.3 Massive $p$-forms on $S^{D}$
and $d^{D}_{n,(p)}$ is the degeneracy (C.2) for the $n$-th transverse $p$-form spherical harmonics on $S^{D}$. Here we focus on the case of the principal series (C.1); the complementary case is obtained by analytic continuation. To derive (C.22) from (4.54), one decomposes the $p$-form into a transverse and longitudinal part: $A_{\mu_{1}\cdots\mu_{p}}=A^{T}_{\mu_{1}\cdots\mu_{p}}+p\nabla_{[\mu_{1}}C_{\mu_{2}\cdots\mu_{p}]}$. The integration over the transverse part $A^{T}$ leads to the $n\geq 1$ terms in (C.22). On the other hand, $C_{\mu_{1}\cdots\mu_{p-1}}$ only enters the action through the mass term in the action (4.54); upon integration, this leads to an infinite product of $m^{2}$. Absorbing this into a local counterterm in a way consistent with locality leads to the n=pn=-p or n=p(D1)n=p-(D-1) term in the range $0\geq n\geq-\lfloor\frac{D-1}{2}\rfloor$ in (C.22). This is done explicitly for $p=1$ in [103].
@@ C.4 Massless $p$-form gauge fields on $S^{D}$
$\mathcal{Z}^{p}_{\rm PI}\left[S^{D}\right]=\left(\frac{\mu^{p+1}\sqrt{2\pi R^{D}{\rm Vol}\left(S^{D}\right)}}{q}\right)^{(-)^{p}}\left(\det\hskip 0.0pt^{\prime}\frac{\Delta_{0}}{\mu^{2}}\right)^{-\frac{(-)^{p}}{2}}\prod_{k=1}^{p}\left(\det\frac{\Delta^{T}_{k}}{\mu^{2}}\right)^{-\frac{(-)^{p-k}}{2}}\;.$ (C.28)